If my standard equation is y+9=-3(x-2) and my answer was 3x+y=7. Was I correct?
Found 2 solutions by fcabanski, RedemptiveMath:
Answer by fcabanski(1391)

y+9=-3(x-2) (distribute -3) ---> y+9 =-3x + 2 (add 3x to both sides and subtract 9 from both sides.) ---> 3x+y = -7.


For the equation to be in slope/intercept form it's y = -3x -7

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)


Answer by RedemptiveMath(80)

You were off by a little mark. You were on the right path, but here is my work showing the correct answer:

y+9 = -3(x-2) (point-slope form)
y+9 = -3x+6
y = -3x-3 (slope-intercept form)
3x + y = -3 (standard form).

I don't know if you multiplied incorrectly or overlooked the distributive part. I don't know if your course wants you to provide the answer in a particular form, so I have included the three basic forms of linear equations.